OpenStudy (anonymous):

How do you get sin(x+2)/x+2 as x approaches -2 to equal 1? I am missing this step and I don't know how to do it.

OpenStudy (anonymous):

replace \(x+2\) by \(u\) making \[\lim_{x\to-2}x+2=\lim_{u\to 0}u\]

OpenStudy (anonymous):

then you have \[\lim_{u\to 0}\frac{\sin(u)}{u}=1\]